dgreer4:

Suppose a hybrid automobile has an average fuel economy of 60 mi/gal and a comparable nonhybrid averages 25 mi/gal . If the hybrid model costs $5000 more than the nonhybrid and gasoline averages $3.75 per gallon which model would be less expensive over a lifespan of 125,000 miles?

ddog:

Dont hold me to it but I believe the hybrid would be less expensive

surjithayer:

cost of fuel for nonhybrid\[=\frac{ 125,000 }{ 25 }\times 3.75=5000 \times 3.75=18750 $\] hybrid fuel used\[=\frac{ 125,000 }{ 60 }=\frac{ 6250 }{ 3 } ~gal\] cost of hybrid fuel=18750+5,000=23750 $ cost of one gallon hybrid fuel \[=\frac{ 23,750 }{ \frac{ 6250 }{ 3 } }=\frac{ 3 }{ 6250 }\times =11.40 ~$\] non hybrid fuel costs less.

surjithayer:

correction in the end (3/6250) *23750=11.40 $